a minor league baseball team plays 125 games in a season. if the team won 17 more than three times as many games as they lost, how many wins and losses did the team have?
Answer by harpazo(655)


x + 3x + 17 = 125
4x + 17 = 125
4x = 125 - 17
4x = 108
x = 108/4
x = 27
The team lost 27 games.
It won 3(27) + 17 games or 98 games.
Nice question....